The least integral value of ‘a’ for which the graphs y=2ax + 1 and y=(a-6)x2 – 2 do not intersect is___.

Grade:11

1 Answers

Vikas TU
7 years ago
Put y = 2ax +1 in y=(a-6)x2 – 2
we get,
(a-6)x2 – 2 – 2ax – 1 =0
(a-6)x^2 – 2ax – 3 = 0
Its D
.i.e roots for x values should be imaginary.